If a computer bill cost 225 dollars
and the Parts cost 75 dollars and 50 dollars for labor. How many hours did it take?

Answer by jim_thompson5910(35256)

Assuming 50 dollars per hour for labor

Let x = # of labor hours

75 + 50x = 225

50x = 225 - 75

50x = 150

x = 150/50

x = 3

It took 3 hours of labor.